stefanie and judy went to the rodeo this weekend. stefanie bought 2 pretzels and 3 cokes for $8.25. Judy bought 3 pretzels and 1 coke for $4.50. how much sis one coke cost?

Answer by checkley77(12569)

2P+3C=8.25
3P+C=4.50 MULTIPLY BY -3 & ADD.
-9P-3C=-13.50
2P+3C=8.25
---------------
-7P=-5.25
P=-5.25/-7/
P=.75 COST OF ONE PRETZEL.
2*.75+3C=8.25
1.50+3C=8.25
3C=8.25-1.50
3C=6.75
C=6.75/3
C=2.25 COST OF ONE COKE.
PROOF:
3*.75+2.25=4.50
2.25+2.25=4.50
4.50=4.50